OpenStudy (lilcathyx):

help, ill give you a medal, and if you can explain it that would be wonderful ! An angle measure of 234 degrees is equivalent to ____ radians. Round your answer to the nearest hundredth when necessary.

OpenStudy (lilcathyx):

dont i have to use the equation\[2\pi(x) \over 360 \]

OpenStudy (lilcathyx):

@pooja195 @haleyelizabeth2017 :O ?

OpenStudy (anonymous):

I will help u if I can

OpenStudy (haleyelizabeth2017):

I'm not sure...sorry!

OpenStudy (anonymous):

im not sure either about this one.

OpenStudy (lilcathyx):

@saynabahmed13 please :) & its okkay @strawberryswing @haleyelizabeth2017

OpenStudy (anonymous):

Well, you are correct. basically by doing \(\frac{\text{angle}^\circ}{360^\circ}\) you get the factor of 'how much of the circle this angle takes'. for example \(180^\circ\) will end up in \(\frac{1}{2}\) and \(360^\circ\) will end up in \(1\). Now by saying that a full circle would be \(2\pi\) radians (because the radius of a circle fits exactly \(2\pi\) times in the perimeter you can multiply it by the factor above and find out how many times the radius fits in the part of circle the angle describes. $$ \frac{\text{angle}^\circ}{360^\circ} \cdot 2\pi = \frac{\pi \cdot \text{angle}^\circ}{180^\circ} $$

OpenStudy (lilcathyx):

234/ 360 x 2pi = pi x 234 / 180 ?

OpenStudy (anonymous):

yes, calculate 234/180, then multiply by pi and you're good =)

OpenStudy (lilcathyx):

234/180 = 1.3.....1.3 x 3.14 =4.082..

OpenStudy (lilcathyx):

4.082 is the answer ?